# Underage Detection through a Multi-Task and MultiAge Approach for Screening Minors in Unconstrained Imagery

### v1

@@ 5.4.2 Effect of architecture and loss tuning on detection and estimation
Above we have shown that the best results for age estimation and underage detection in the wild are obtained with a FaRL + MLP model with four under-age detectors, trained with the cleaned overall age dataset (without CACD) in an age-equalized manner. However, there are several parameters that can be investigated further. Here, we investigate two variation points, namely the MLP architecture and three parameters of the loss function, the balancing parameter $\alpha$, the focusing parameter $\gamma$, and the age gap. Table 8 collects the performance measures of a collection of models that vary these variation points.
Test ASORES-39k (restricted test) ASWIFT-20k (wild test)
F2 F2 F2 F2 RMSE F2 F2 F2 F2 RMSE
MLP $G$ $\gamma$ $\alpha$ $T\!=\!12$ $T\!=\!15$ $T\!=\!18$ $T\!=\!21$ (years) $T\!=\!12$ $T\!=\!15$ $T\!=\!18$ $T\!=\!21$ (years)
WS 0 0 0.959 0.938 0.875 0.875 5.602 0.933 0.909 0.813 0.851 6.497 (D)
0 0 yes 0.961 0.915 0.869 0.878 5.635 0.935 0.879 0.804 0.850 6.630
0 2 0.954 0.922 0.867 0.872 5.619 0.905 0.876 0.805 0.848 6.470
1 0 0.957 0.954 0.879 0.877 5.582 0.931 0.911 0.810 0.852 6.499
2 0 0.958 0.947 0.873 0.873 5.588 0.944 0.903 0.778 0.854 6.469
3 0 0.949 0.934 0.854 0.874 5.591 0.917 0.884 0.819 0.847 6.441
rel 0 0.960 0.946 0.856 0.873 5.557 0.948 0.911 0.822 0.846 6.408 (E)
I-512 0 0 0.962 0.941 0.876 0.883 5.626 0.946 0.913 0.801 0.852 6.626
I-256 0 0 0.965 0.948 0.869 0.878 5.639 0.946 0.918 0.809 0.853 6.548
3 0 0.961 0.931 0.864 0.872 5.647 0.947 0.881 0.840 0.851 6.457
3 2 0.931 0.921 0.827 0.876 5.644 0.904 0.884 0.810 0.848 6.533
rel 0 0.965 0.951 0.857 0.875 5.656 0.955 0.916 0.833 0.854 6.458 (F)
I-128 0 0 0.958 0.937 0.868 0.866 5.670 0.920 0.904 0.800 0.845 6.489
@@
Table 8: Performance measures of underage detection at age thresholds $12,15,18,21$, and age regression. Common parameters: heads = age + 12,15,18,21, training without CACD and with resampling (like model D in Table 7). Variation points are the architecture of the MLP [weight-sharing MLP (WS) vs. = two-layer independent MLP with intermediate width X (I-X), see Sec. 4.1], the age gap $G$ (“rel” means age gap according to Eq. 9), the focusing parameter $\gamma$, and the class reweighting $\alpha$. Top three values in bold face.
As seen already in Figure 7 and Table 5 for under-18 discrimination, the alpha parameter does not improve accuracy for the other age thresholds either, given that resampling is already in place.
@@
Two experiments with the gamma parameter increased to $\gamma=2$ are shown (3rd vs. 2nd line in the WS-block; 3rd line vs. 2nd line in the I-256 block in Table 8), where the focal loss (Eq. 5) focuses on the most “difficult” patterns. This leads to no clear improvement either. Our hypothesis for this failure is that focusing on the difficult examples may lead to overfitting to edge cases and label noise.
@@
The next variation point, introducing an age gap, is motivated by the opposite approach, namely by removing edge cases from the loss function (Section 4.3). This adds the configuration of each of the discriminators to the configuration space. We have tried fixed age gaps of ($G=1,2,3$, i.e. with age gap $[-G,G]$ with respect to the age threshold) years for all four discriminators. The data for MLP = WS and $G=0,1,2,3$ in Table 8 does not show clear trends on ASORES-39k, but it does on ASWIFT-20k: Both underage detection and age regression benefit from the age gap, the F2 of $T=12$ increases from 0.933 to 0.944 at $G=2$; the F2 of $T=18$ increases from 0.813 to 0.819 at $G=3$. Since the optimal value of $G$ turns out to vary as a function of the age threshold, we propose a relative scaling of the age gap scales with the age threshold:
age $\textrm{age gap}=\begin{cases}12+[-2,+2],&T=12,\\ 15+[-2,+3],&T=15,\\ 18+[-3,+3],&T=18,\\ 21+[-3,+4],&T=21.\end{cases}$ (9)
Due to the integer nature of the age labels, the age gaps of odd width had to be chosen asymmetrically. This configuration leads to the model labeled as (E) in Table 8. It shows improvements over model D for under-12, under-15 detection and age regression, but not in under-18 and under-21 detection.
@@
In the lower part of Table 8, we show the accuracy measures of models where this weight-sharing two-layer MLP is replaced with a regular MLP of two (independent) linear layers, each followed by a ReLU activation function. Input and output dimensions are kept at 512, but the intermediate width is varied to adjust the capacity of the MLP. This “bottleneck” is to make up for the regularizing effect of the shared weights in the original SW implementation. The results show that the MLP with intermediate bottleneck (I-256) achieves slightly better under-age detection F2 scores, but are worse in age estimation. If combined with the gap of Eq. 9 (gap = rel), however, the resulting model F improves consistently over model D on the ASWIFT-20k test.
@@
In conclusion, in this Subsection 5.4.2, we have shown that the focal-loss parameters $\alpha$ and $\gamma$ have only minor effects and cannot improve the performance of model D. The gap parameter, in turn, has the potential for some improvements if configured correctly. Finally, we have shown that the MLP of independent linear layers and a bottleneck of width 256 can replace the weight-sharing MLP reaching comparable accuracy. Combined with the age gap (model F), it consistently improves over model D in the ASWIFT-20k wild test, albeit with slight degradation in the ASORES-39k test for under-18 detection and age regression. Still, such a model with more even accuracy across domains is desirable.
@@ 6 Conclusions
In this work, we propose a multitask architecture for facial age estimation and underage detection that is robust in unconstrained imagery. The model performs age regression and binary underage classification at four critical legal thresholds (12, 15, 18, and 21 years) simultaneously.
@@
The model is trained on on a curated multi-source dataset excluding the noisy CACD2000 dataset and by removing label noise. To address the imbalance in age distribution we propose a resampling strategy where the under-represented age groups (minors) are sampled more frequently in the training than others. This approach is superior to just adding weight factors ($\alpha$) to the loss function. Removing edge cases by introducing an “age gap” in the binary loss functions improves the generalization of the trained model.
For evaluation, we designed a challenging benchmark, the ASWIFT-20k wild test, which deliberately includes images with extreme head poses, facial expressions and illumination conditions.
The multitask approach of combining the age-regression head of the model with several binary discriminators for non-adult detection turned out to be beneficial both for the age regression task and for the non-adult discrimination tasks. The reason for this is the synergies between age estimation and the different underage discriminators. The tasks are inherently related (knowing the exact age solves both), but the error measures and the loss function are different.
The present work has established a new benchmark for facial age estimation and under-age detection in the wild, but it also opens the door for several future research lines, both both in methodology and dataset development.
-Methodology: Super-resolution approaches for bad imaging conditions
Wild data often suffers from low resolution, poor lighting, and occlusions, which hinder age estimation accuracy. Here, an effective approach could involve the use of super-resolution techniques to enhance the quality of facial inputs. Recent advances in deep learning have made it possible to generate realistic, high-detail reconstructions using adversarial networks, as in ESRGAN [33] or Real-ESRGAN [34]. However, it is not clear a priory whether visual improvements lead to better predictions or introduce bias (e.g., by smoothing wrinkles).
-Demographic aspects of the data
Improving the training set with more age-labeled images, especially for underrepresented combinations of age, ethnicity, gender, and image conditions, is crucial. Possibly generative methods for facial-expression transfer [35] can help to fill some the gaps in the training data.
-Benchmark Augmentation
The age models have to be tested on genuine data (no generative methods) in order to evaluate the accuracy. Such a test must also prove whether synthetic data used in the training helps to solve the genuine task or not.
The Wild Test is what could be achieved with the currently available data. It would be desirable, however, to better control the demographic distribution across the different sub-tests. Currently, for example, strong head poses come mostly from the Dartmouth Database of Children’s Faces, i.e. age and head pose are correlated. Ideally, the Wild Test could be split into gender-and-ethnicity groups and every group would have a similar distribution in age and image conditions (head pose, facial expression, illumination). Then a quantitative evaluation of the fairness of facial age algorithms with respect to demographic groups could be made and identify any potential biases.
Such an equalization of the test data could be improved by sub-sampling over-represented data, by obtaining more under-represented data, and by upsampling under-represented data (possibly including some augmentation methods, like reflection or reducing the resolution, but no uncontrolled generative methods).

### v2

@@ 5.4.2 Effect of architecture and loss tuning on detection and estimation
Above, we have shown that the best results for age estimation and underage detection in the wild are achieved with a FaRL + MLP model with four underage detectors, trained on the cleaned overall age dataset (without CACD) in an age-equalized manner. However, there are several further variation points to be investigated, namely the MLP architecture and three loss-function parameters: the class balancing parameter $\alpha$, the focusing parameter $\gamma$, and the age gap $G$. Table 7 collects the performance measures of a collection of models that vary these factors.
ASORES-39k (restricted test) ASWIFT-20k (unconstrained)
F2 F2 F2 F2 MAE F2 F2 F2 F2 MAE
MLP $G$ $\gamma$ $\alpha$ T=12T\!=\!12 T=15T\!=\!15 T=18T\!=\!18 T=21T\!=\!21 (years) T=12T\!=\!12 T=15T\!=\!15 T=18T\!=\!18 T=21T\!=\!21 (years)
WS 0 0 0.959 0.938 0.875 0.875 4.062 0.933 0.909 0.813 0.851 4.468 (D)
0 0 yes 0.961 0.915 0.869 0.878 4.094 0.935 0.879 0.804 0.850 4.631
0 2 0.954 0.922 0.867 0.872 4.083 0.905 0.876 0.805 0.848 4.498
1 0 0.957 0.954 0.879 0.877 4.044 0.931 0.911 0.810 0.852 4.473
2 0 0.958 0.947 0.873 0.873 4.060 0.944 0.903 0.778 0.854 4.491
3 0 0.949 0.934 0.854 0.874 4.063 0.917 0.884 0.819 0.847 4.488
rel 0 0.960 0.946 0.856 0.873 4.034 0.948 0.911 0.822 0.846 4.450 (E)
I-512 0 0 0.962 0.941 0.876 0.883 4.041 0.946 0.913 0.801 0.852 4.572
I-256 0 0 0.965 0.948 0.869 0.878 4.065 0.946 0.918 0.809 0.853 4.520
3 0 0.961 0.931 0.864 0.872 4.081 0.947 0.881 0.840 0.851 4.496
3 2 0.931 0.921 0.827 0.876 4.075 0.904 0.884 0.810 0.848 4.556
rel 0 0.965 0.951 0.857 0.875 4.068 0.955 0.916 0.833 0.854 4.475 (F)
I-128 0 0 0.958 0.937 0.868 0.866 4.075 0.920 0.904 0.800 0.845 4.504
@@
Table 7: Performance measures of underage detection at age thresholds $12,15,18,21$, and age regression. Common parameters: heads = age + 12,15,18,21, training without CACD and with resampling (like model D in Table 6). Variation points are the architecture of the MLP [weight-sharing MLP (WS) vs. two-layer independent MLP with intermediate width $m$ (I-$m$), see Sec. 4.1], the age gap $G$ (“rel” means age gap according to Eq. 9), the focusing parameter $\gamma$, and the class reweighting $\alpha$. Top three values in bold face.
@@
As seen in Figure 7 and Table 5 for under-18 discrimination, the $\alpha$ parameter does not improve accuracy for the other age thresholds either, if resampling is in place. Two experiments with the gamma parameter increased to $\gamma=2$ are shown (3rd vs. 2nd line in the WS-block; 3rd line vs. 2nd line in the I-256 block in Table 7), where the focal loss (Eq. 5) focuses on the most “difficult” patterns. This leads to no clear improvement either. Our hypothesis for this failure is that focusing on the difficult examples may lead to overfitting to edge cases and label noise.
@@
The next variation point, introducing an age gap, is motivated by the opposite approach, namely by removing edge cases from the loss function (Section 4.3). This adds the configuration of each discriminator to the configuration space. We have tried fixed age gaps of ($G=1,2,3$, i.e. with age gap [G,G][-G,G] with respect to the age threshold) for all four discriminators. The data for MLP = WS and $G=0,1,2,3$ in Table 7 does not show clear trends on ASORES-39k, but it does on ASWIFT-20k: Underage detection benefits from the age gap, but not at the same gap parameter: the F2 of $T=12$ increases from 0.933 to 0.944 at $G=2$; the F2 of $T=18$ increases from 0.813 to 0.819 at $G=3$. Thus, the optimal value of $G$ seems to scale with the age threshold $T$, as already motivated in Section 4.3, Eq. 9. This configuration leads to the model labeled as (E) in Table 7. It shows improvements over model D for under-12, under-15 detection and age regression, but not in under-18 and under-21 detection.
@@
In the lower part of Table 7, we show the results of models where this weight-sharing two-layer MLP is replaced with a regular MLP of two (independent) linear layers, each followed by a ReLU activation function. Input and output dimensions are kept at 512, but the intermediate width is varied to adjust the capacity of the MLP. This “bottleneck” is to make up for the regularizing effect of the shared weights in the original SW implementation. The results show that the MLP with an intermediate bottleneck (I-256) achieves slightly better underage detection F2 scores, but is worse in age estimation. If combined with the gap of Eq. 9 ($G$ = rel), however, the resulting model F improves consistently over model D on the ASWIFT-20k test.
@@
In this Subsection 5.4.2, we have shown that the focal-loss parameters $\alpha$ and $\gamma$ have only minor effects and cannot improve the performance of model D. The gap parameter, in turn, has the potential to improve it, if configured correctly. Finally, we have shown that the MLP of independent linear layers and a bottleneck of width 256 can replace the weight-sharing MLP, reaching comparable accuracy. Combined with the age gap (model F), it consistently outperforms model D in the ASWIFT-20k wild test, albeit with slight degradation in the ASORES-39k test for under-18 detection and age regression. Still, a model with more even accuracy across domains is desirable.
5.5 Ablation study and comparison with baselines
After designing age estimation and minor detection models based on the frozen FaRL backbone, in particular the model F listed in Table 7, three questions remain to be answered. (a) Is the frozen backbone the best choice, or can further accuracy be gained by fine-tuning? (b) Given the composed training set and the resampling scheme, how important is the architecture of the age estimator and the minor-detection heads? (c) How does our model perform compared to state-of-the-art methods?
5.5.1 Age estimation
Table 8 shows the mean absolute regression error of several experiments addressing these questions. The first line, Age + 4 $\times$ Minor, no fine-tuning, is our Model F from above.
ASORES ASORES ASWIFT ASWIFT
Prediction heads Fine-tuning 39k OOD 18k 20k OOD 10k
$\mathrm{Age}+4\times\mathrm{minor}$ no fine-tuning 4.068 4.212 4.475 4.807 (F)
from 10th block 3.925 4.146 4.527 5.038
from 5th block 3.946 4.220 4.586 5.204
full backbone 3.990 4.220 4.628 5.255
Age only no fine-tuning 4.054 4.228 4.460 4.805
from 10th block 4.011 4.266 4.642 5.292
from 5th block 4.069 4.562 4.638 5.440
Direct regression no fine-tuning 4.049 4.082 4.483 4.703
from 10th block 4.005 4.254 4.461 4.886
from 5th block 4.169 4.845 4.733 5.730
DEX: VGG-16, IMDB-WIKI. [Rothe_2015_DEX] 6.577 7.844 7.632 9.323
MiVOLO, vision transformer, [Kuprashevich_2024_MiVOLO] 4.560 4.355 5.653 5.738
Table 8: Fine-tuning and comparison to two baseline methods (MAE in years). The out-of-distribution (OOD) variants of the tests are restricted to the datasets CACD2000, FG-NET, CASIA-Face-Africa, and Dartmouth, i.e., those not used in the overall training set, and images where the MiVOLO face detector failed were excluded, allowing for a fair comparison with the external baseline models DEX and MiVOLO.
We compare the FaRL backbone with three different setups for the prediction heads (age + 4 $\times$ minor detection; age only; direct age regression) and further vary the degree of fine-tuning of the backbone (no fine-tuning; fine-tuning from the 10th resblock on; fine-tuning from the 5th resblock on; fine-tuning of the entire vision transformer). The impact of these changes on the prediction accuracy depends on the test case: Sometimes, a bit of fine-tuning helps, but not always: For the ASORES test, moderate fine-tuning (from the 10th resblock onward) further reduces the estimation error in all three setups. However, ASWIFT and ASWIFT-OOD tests are best without fine-tuning. The differences between the losses (age + 4 $\times$ minor detection, age only, and direct regression) are not significant. Since the main motivation of our paper is robust performance under adverse conditions, we prioritize the ASWIFT test. The conclusion is that neither fine-tuning nor changes to the loss function can significantly improve age-regression accuracy, at least with the current training data.
The last block of the table shows the mean absolute errors achieved by the classic DEX model [Rothe_2015_DEX] and the state-of-the-art MiVOLO model [Kuprashevich_2024_MiVOLO]. For a fair comparison, consider the columns ASORES-OOD and ASWIFT-OOD. Our model “F” outperforms both baseline models. While the MiVOLO model is close to ours on the ASORES-OOD test, our model has a significant advantage on the ASWIFT-OOD test.
Multi-task vs. single-task
To evaluate the synergy of the tasks of age estimation and underage detection, we compare the model (F) to a single-task underage detector with the same parameters. If the age gap is preserved, the single-task model fails: the separation of adults and minors is so bad that no recall better than 0.92 can be reached. When the age gap is removed, this failure is healed, and the model reaches reasonable accuracy again, with a performance similar to Model F on ASORES-39k. On the ASWIFT-20k test, however, we measure Pr = 0.439, F2 = 0.791, which is worse than the multitask model F (Pr = 0.512, F2 = 0.833, see also Table 7). Thus, the age gap works well only in the multitask setup, last hidden layers are shared with the age regressor, but then, this setup leads to enhanced generalization.
5.5.2 Underage detection and age classification with partially occluded faces
Model MAE Classification
Chaves et al. [Chavez_VISAPP2020], (IMDB, 4000 img/age, only original) 5.02 y Acc = 33.9%
Chaves et al. [Chavez_VISAPP2020], (IMDB, 4000 img/age, original + occluded) 4.07 y Acc = 39.2%
Model F (ours) 4.11 y Acc = 47.6%
Table 9: Performance comparison with Cháves et al. on the occluded + original test of Ref. [Chavez_VISAPP2020].
We have evaluated our model F on the test set of [Chavez_VISAPP2020], which is an age-balanced dataset ranging from 0 to 25 years with 1000 images per year and composed of images from the Diversity in Faces dataset [Merler_2019_DIF] (86%), IMDB-WIKI [Rothe_2018_IMDBWIKI] (5.2%), FG-NET [Lanitis_2002_FGNET] (0.01%), UTKFace [Zhang_2017_UTKFace] (5.8%), APPA-REAL [Agustsson_2017_AppaReal] (2%), and AgeDB [Moschoglou_2017_AgeDB] (1.3%). Only the last three datasets (9.1%) have a potential overlap with our training set since the exact train-test splits are not reproducible anymore. Each image is tested twice, with and without a black bar drawn as an artificial occlusion over the eye region. The figures of merit are the MAE measuring the regression error and the mean accuracy for classifying faces into five age groups with boundaries 0, 6, 11, 16, 18, and 26 years. In Table 9, we compare our model F to the values reported in Ref. [Chavez_VISAPP2020]. Even though our model is not specifically trained for occluded faces, its age estimation accuracy ($\mathrm{MAE}=4.11\,\mathrm{y}$) comes very close to that of the strongest model of [Chavez_VISAPP2020], trained with original and occluded images ($\mathrm{MAE}=4.07\,\mathrm{y}$). In terms of classification accuracy, model F clearly outperforms all models of Chaves et al. [Chavez_VISAPP2020].
5.6 Analysis of demographic effects and failure modes
In this section, we analyze in more detail how our model F responds to shifts in the data distribution and under what conditions it is likely to fail.
Figure 9: Error distribution by age and yaw angle on the ASWIFT-20k test.
In Figure 9 we show the distribution of the age estimation error, divided into groups of age and yaw angle. Within each age group, the MAE and the distribution of the error vary only very little with the yaw angle, highlighting the robustness of the underlying FaRL backbone. We do observe, however, the well-known pattern that age estimation in terms of absolute regression errors is easier at young ages than at older ages.
Under-12 Detection Under-15 Detection Under-18Detection Under-21 Detection
Test Pr F1 Re Pr F1 Re Pr F1 Re Pr F1 Re MAE
Source: ASORES-39k 0.88 0.93 0.99 0.82 0.90 0.99 0.56 0.71 0.99 0.60 0.75 0.99 4.07
Source: ASWIFT-20k 0.83 0.90 0.99 0.73 0.83 0.98 0.51 0.67 0.99 0.56 0.71 0.98 4.48
Gender: Male 0.90 0.94 0.99 0.81 0.89 0.99 0.55 0.70 0.98 0.56 0.71 0.98 3.94
Gender: Female 0.80 0.89 0.99 0.70 0.81 0.98 0.45 0.62 0.99 0.53 0.69 0.99 4.71
Race: Asian 0.89 0.94 1.00 0.63 0.77 0.99 0.26 0.40 0.94 0.36 0.53 0.99 3.88
Race: White 0.87 0.93 0.99 0.85 0.91 0.99 0.61 0.76 1.00 0.68 0.80 0.99 4.17
Race: Black 0.29 0.44 0.95 0.14 0.24 0.90 0.24 0.38 0.94 0.42 0.59 0.97 4.50
Pose: Extreme Yaw 0.86 0.93 1.00 0.93 0.96 0.99 0.92 0.96 1.00 0.92 0.96 1.00 2.92
Pose: Extreme Pitch 0.88 0.93 1.00 0.74 0.85 0.99 0.47 0.64 0.99 0.56 0.71 0.99 4.47
Image: Dark 0.77 0.86 0.98 0.67 0.80 0.97 0.49 0.65 0.98 0.52 0.68 0.98 4.36
Image: Blurry 0.78 0.87 0.98 0.52 0.68 0.98 0.31 0.47 0.97 0.39 0.55 0.97 4.66
Image: Low Contrast 0.74 0.85 0.98 0.43 0.59 0.96 0.28 0.44 0.97 0.43 0.59 0.96 4.70
Image: Low Saturation 0.68 0.80 0.96 0.32 0.47 0.96 0.21 0.35 0.93 0.40 0.57 0.98 4.41
Image: High Saturation 0.72 0.83 0.99 0.64 0.77 0.96 0.49 0.65 0.99 0.54 0.70 0.99 4.62
Expr: High Arousal 0.89 0.94 0.99 0.85 0.92 0.99 0.65 0.78 0.99 0.67 0.80 0.99 4.11
Expr: Low Valence 0.83 0.90 0.99 0.82 0.90 0.99 0.60 0.75 0.99 0.63 0.77 0.99 4.21
Expr: High Valence 0.89 0.94 0.99 0.79 0.88 0.99 0.49 0.65 0.99 0.57 0.72 0.98 4.50
DB: CASIA-Face-Africa 0.11 0.19 0.89 0.10 0.18 0.81 0.21 0.34 0.98 0.50 0.66 0.94 7.22
DB: FG-NET 0.86 0.91 0.96 0.79 0.88 0.99 0.78 0.88 0.99 0.88 0.93 0.98 3.63
Table 10: Performance measures of Model F across demographic and image-condition subsets. The top rows show the ASORES-39k and ASWIFT-20k tests (see also the rows marked with (F) in Tables 7 and 8). Further rows show subsets of the union of ASORES-39k and ASWIFT-20k. Gender and Race: estimates of the FairFace model [Karkkainen_2021_FairFace]; Pose: highest 10% of the InsightFace angles; Image properties: highest/lowest 20% of the image properties; Expression: lowest/highest 10% of the arousal and valence parameter of [Vonikakis_2021_expression] (details Sec. 3.2.3). The last rows show the CASIA-Face-Africa subset (3375 images) of ASWIFT-20k and the FG-NET dataset (1k images). In each column, the four worst-performing cases are highlighted in italics.
A more extensive analysis is given in Table 10, which lists performance measures, both for underage detection and for age regression, for different subsets of the union of ASORES-39k and ASWIFT-20k. In most cases, the recall (Re) remains high and the precision (Pr) does not drop too much. In particular, there are no problems with large pose angles and strong expressions. However, we do observe relevant deterioration of performance for the Female group (age regression), Asian and Black groups (problems with false adult detections, but recall stays high), and the Blurry, Low Contrast, and Low Saturation groups (problems with both underage detection and age regression).
These observations should be interpreted cautiously, as the test data is unevenly distributed. Females are strongly represented in CACD2000, while males are overrepresented in MORPH. Asians mostly come from the AFAD dataset, and Blacks from MORPH and CASIA-Face-Africa, both of lower image quality.
The last two rows show two cases of cross-dataset validation, i.e., data from datasets that were not used in the training at all, not just hold-out sets. The row CASIA-Face-Africa shows results of the 3375 lowest quality webcam images from the CASIA-Face-Africa dataset selected for the ASWIFT-20k test (see Table 1). In this doubly difficult case, the model reaches its limits; the precision breaks down, but the recall remains high. The results measured with the FG-NET dataset show that the MAE of $3.63$ years is better than the values in Tables 2 and 4 and values reported elsewhere (Paplhám & Franc [Paplham_CVPR2024_benchmark] report $4.41$ years for a pre-trained regression model trained on UTKFace).
@@ 6 Conclusions
In this work, we propose a multitask architecture for facial age estimation and underage detection that is robust in unconstrained imagery. The model performs age regression and binary underage classification at four critical legal thresholds (12, 15, 18, and 21 years) simultaneously. The model leverages the face representation of the FaRL backbone and is trained on a curated multi-source dataset, excluding the noisy CACD2000 dataset, and by removing label noise. To address the imbalance in age distribution, we propose a resampling strategy in which underrepresented age groups (minors) are sampled more frequently in the training set than the others. This approach is superior to adding weight factors ($\alpha$) in the loss function. In conjunction with the age-estimation task in the multi-task architecture, removing edge cases from the binary loss functions by introducing an “age gap” improves the generalization of the trained model. Thus, the multi-task setup is able to exploit synergies between the related tasks, age-regression and underage detection.
For evaluation, we designed a challenging benchmark, the ASWIFT-20k wild test, which deliberately includes images with extreme head poses, facial expressions, and illumination conditions. We have analyzed our model’s performance on subsets of the test data, defined by head pose, facial expression, demographic groups, and imaging conditions. On the plus side, our model is robust to pose and expression variation and shows lower estimation errors for young subjects than for older ones. There are weaknesses, however, for Black subjects and under difficult imaging conditions.
More investigation is needed to identify and remedy such biases. Expanding the training set with additional age-labeled images for underrepresented combinations of age, ethnicity, gender, and image conditions could be essential. Generative data augmentation techniques (e.g., facial expression transfer) may address demographic gaps while maintaining diversity and realism. Low-quality images could be enhanced by applying super-resolution techniques that preserve the age-relevant features of the image. Additionally, we intend to refine the ASWIFT-20k benchmark for a more balanced distribution, reducing correlations between parameters like pose, race, or age, allowing a more rigorous and fair evaluation of model robustness.
Finally, we acknowledge the ethical considerations related to the use of facial analysis technology. While our research is motivated by its application to minor protection, specifically to CSEM detection, we recognize that age estimation tools could be misused. Scenarios such as pervasive surveillance and automated discrimination based on age are significant concerns. Therefore, we stress the responsible deployment of these systems. Their use should be limited to specific, ethically evaluated, and legally sanctioned cases. Moreover, given the identified bias and performance gaps, we emphasize the importance of human supervision in any critical deployment to grant fairness, correct errors and prevent harm.
The code of this work is released at https://gitlab.com/gvis-lab/MultiAge.